# TI-DeepONet: Learnable Time Integration for Stable Long-Term Extrapolation

### v2

@@ Abstract
Accurate temporal extrapolation presents a fundamental challenge for neural operators in modeling dynamical systems, where reliable predictions must extend significantly beyond the training time horizon. Conventional deep operator network (DeepONet) approaches rely on two inherently limited training paradigms: fixed-horizon rollouts, which predict complete spatiotemporal solutions while disregarding temporal causality, and autoregressive formulations, which accumulate errors through sequential predictions in time. We introduce TI-DeepONet, a framework that integrates neural operators with adaptive numerical time-stepping techniques to preserve the underlying Markovian structure of dynamical systems while substantially mitigating error propagation in extended temporal forecasting. Our approach reformulates the learning objective from direct state prediction to the approximation of instantaneous time-derivative fields, which are subsequently integrated using established numerical schemes. This design naturally supports continuous-time prediction and enables the use of higher-precision integrators at inference than those employed during training, striking a balance between computational efficiency and predictive accuracy. We further develop TI(L)-DeepONet, which incorporates learnable coefficients for intermediate slopes in a multi-stage numerical integration scheme, thereby adapting to solution-specific variations and enhancing predictive fidelity. Rigorous evaluation across four canonical partial differential equations (PDEs) spanning diverse, high-dimensional, chaotic, dissipative, and dispersive dynamics shows that TI(L)-DeepONet marginally outperforms TI-DeepONet, with both frameworks achieving significant reductions in relative $L_{2}$ extrapolation error: approximately 81% compared to autoregressive implementations and 70% compared to fixed-horizon approaches. Notably, both methods maintain stable predictions over temporal domains extending to nearly twice the training interval. This research establishes a physics-aware operator learning paradigm that bridges neural approximation with numerical analysis principles, preserving the causal structure of dynamical systems while addressing a critical gap in long-term forecasting of complex physical phenomena.
@@ 1 Introduction
Time-dependent partial differential equations (PDEs) are typically solved using spatial discretization (e.g., finite element or volume) combined with time-stepping schemes such as Runge–Kutta or Adams–Bashforth. While accurate, such solvers can be prohibitively expensive for high-dimensional or many-query problems. Explicit integrators are constrained by stability conditions (e.g., CFL), while implicit schemes require solving large systems of equations at each step. Moreover, classical solvers generalize poorly: each new initial or boundary condition requires a fresh simulation.
@@
Neural operators (NOs) offer an attractive alternative by learning mappings from input functions (e.g., initial conditions) to solutions, bypassing discretization and enabling fast inference. Models like DeepONet [lu2021learning] and FNO [lifourier] have shown promise in parametric PDE settings. However, extending NOs to time-evolving systems introduces new challenges. The standard approach is to train a model that maps the solution at time $t$ to the solution at time $t+\Delta t$, and roll this forward autoregressively (see Figure 1). Unfortunately, this approach suffers from severe error accumulation, as each prediction feeds into the next, compounding inaccuracies and often leading to instability. Alternatively, one can train a model to predict an entire trajectory (“full rollout”), which avoids stepwise feedback but tends to degrade in accuracy when extrapolating beyond the training time window. Crucially, most existing works [lin2023learning, lifourier, wang2023long, lu2021learning] only evaluate models within the training time span, neglecting the need for reliable extrapolation, a capability that is essential for real-time engineering applications and control tasks. Moreover, existing work [xu2023transfer, diab2024temporal, liu2022deeppropnet] on autoregressive and full rollout operator models tends to impose a discrete-time Markovian assumption, which misaligns with the underlying continuous-time dynamics of PDEs. These limitations often result in unstable or inaccurate long-term predictions. Thus, a key barrier remains: NOs struggle not only with long-horizon accuracy but also with extrapolating into unseen temporal domains, a prerequisite for trustworthy deployment.
@@
Several recent works attempt to address these limitations. For example, Diab et al. [diab2025temporal] augment DeepONet with an additional temporal branch to capture dependencies, while Michalowska et al. [michalowska2024neural] employ recurrent networks (RNNs, LSTMs, GRUs) atop neural operator outputs to better handle sequences. Other approaches introduce memory modules [buitragobenefits, he2024sequential, hu5149007deepomamba] to compensate for the non-Markovian nature of rollout-based methods. However, these strategies largely rely on generic time-series architectures rather than tackling the issue from a dynamical systems perspective.
@@
In this work, we introduce a novel approach inspired by classical numerical analysis. Our proposed TI-DeepONet framework embeds time integration directly into the operator learning process for dynamical systems. Rather than predicting future solution states directly, we train a DeepONet to approximate the instantaneous time derivative $\partial u/\partial t$. During training, this derivative operator is coupled with a Runge-Kutta numerical integrator to evolve the state forward in time. The predicted future state is then compared against the ground truth in the loss function, enabling the network to learn dynamics that respect temporal causality and numerical stability. At inference time, we leverage high-order multistep integrators (such as Adams-Bashforth/Adams-Moulton predictor-corrector schemes) with refined timesteps to reliably propagate solutions based on the learned time derivative operator. This operator learning framework which maps between function spaces rather than from function spaces to vector spaces as in standard neural networks and neural ODEs [chen2018neural] - provides the flexibility to modify time-stepping parameters between training and testing phases, thereby ensuring stable predictions. We further introduce TI(L)-DeepONet, an enhanced variant that incorporates learnable weighting coefficients conditioned on the system state at each timestep. This adaptation dynamically adjusts the integration process based on local solution characteristics.
@@
We demonstrate the effectiveness of our approach on four benchmark PDEs: (1) 1D Burgers’, (2) 1D Korteweg-de Vries (KdV), (3) 1D Kuramuto-Sivashinsky (KS), and (4) 2D inviscid Burgers’ equations. We compare our methods against four standard neural operator (NO) baselines: (1) autoregressive DeepONet (DON AR), (2) full rollout DeepONet (DON FR), (3) autoregressive FNO (FNO AR), and (4) full rollout FNO (FNO FR). Across all cases, TI-DeepONet and its adaptive variant TI(L)-DeepONet significantly outperform standard autoregressive and full rollout operator models, achieving stable long-term predictions and accurate extrapolation for time horizons approximately twice as long as the training interval.
@@ 2.1 Deep operator network (DeepONet)
Deep operator network (DeepONet) are grounded in the universal approximation theorem for nonlinear operators [chen1995universal], and are designed to learn mappings between infinite-dimensional input and output function spaces. A standard DeepONet architecture comprises two subnetworks: a branch network and a trunk network. The branch network encodes the input function $\mathbf{v}(\eta)$, evaluated at a set of fixed sensor locations $\{\eta_{1},\eta_{2},\dots,\eta_{m}\}$, producing a set of coefficient weights. The trunk network takes as input the query coordinates $\zeta=(x,y,z,t)$ and outputs the corresponding basis functions.
@@
In general, operator learning frameworks like DeepONet aim to learn a solution operator that maps any functional input, such as initial conditions (ICs), boundary conditions (BCs), or source terms, to the corresponding spatiotemporal solution field. However, this mapping is inherently non-causal with respect to time: the operator does not explicitly enforce the Markovian nature of dynamical systems, where the future state depends on the present (or past) states. As a result, direct operator prediction over time often fails to preserve temporal coherence, particularly for long-term simulations.
@@ 2.2 Time Integrator embedded deep operator network (TI-DeepONet)
To mitigate these limitations, we propose the TI-DeepONet architecture. Rather than directly predicting the next state $u^{i+1}$, the network learns the right-hand side (RHS) of the time-dependent PDE (Eq. 1), i.e., the time derivative $u_{t}=\mathcal{F}(t,\mathbf{x},u,u_{\mathbf{x}},u_{\mathbf{xx}},\dots)$. This learned derivative is passed to a numerical time integrator to compute the next solution state from the current one. In particular, we use the RK4 scheme to advance the solution in time. For a timestep $\Delta t$, the RK4 update rule is given by, $k_{1}=\mathcal{F}(t^{i},u^{i})$, $k_{2}=\mathcal{F}\left(t^{i}+\tfrac{\Delta t}{2},u^{i}+\tfrac{\Delta t}{2}k_{1}\right)$, $k_{3}=\mathcal{F}\left(t^{i}+\tfrac{\Delta t}{2},u^{i}+\tfrac{\Delta t}{2}k_{2}\right)$, $k_{4}=\mathcal{F}\left(t^{i}+\Delta t,u^{i}+\Delta t\cdot k_{3}\right)$, and $u^{i+1}=u^{i}+\Delta t\left(\tfrac{1}{6}k_{1}+\tfrac{2}{6}k_{2}+\tfrac{2}{6}k_{3}+\tfrac{1}{6}k_{4}\right)$. This formulation leads to an end-to-end differentiable pipeline, where the loss is computed between the predicted solution $\hat{u}^{i+1}$ and the ground truth $u^{i+1}$, and gradients are backpropagated through the time integrator to update the network parameters. However, the RK4 slope coefficients, $\{\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}\}$, are fixed.
@@ 2.3 Learnable Time Integrator embedded deep operator network (TI(L)-DeepONet)
For numerical stability and interpretability, we normalize the weights using softmax activation function, such that: $\tilde{\alpha}_{j}=\frac{exp(\alpha_{j})}{\sum_{\ell=1}^{4}exp(\alpha_{\ell})},\forall j=1,2,3,4$. This adaptive scheme allows the model to adjust the influence of each intermediate slope dynamically, thereby compensating for the approximation error [lanthaler2022error] of the DeepONet prediction, enabling more robust predictions in stiff, nonlinear, or highly transient regimes. For example, in regions with rapid spatial changes, the model may place greater weight on later-stage evaluations like $k_{4}$, while in smoother regions, earlier evaluations may suffice.
@@ 3 Results
To evaluate the efficacy of our framework, we consider four canonical PDEs: (1) 1D Burgers’ equation, (2) 1D KdV equation, (3) 1D KS equation, and (3) 2D inviscid Burgers’ equation. Model performance is assessed using the relative $L_{2}$ error, defined as $\text{error}=\frac{\|u_{\text{pred}}-u_{\text{true}}\|_{2}}{\|u_{\text{true}}\|_{2}}$, where $u_{\text{pred}}$ and $u_{\text{true}}$ denote the predicted and ground truth solutions, respectively. Table 1 and Figure 3 summarize the comparative training and extrapolation accuracy across all methods for all applications. To evaluate the statistical variability of the performance of the model, we report the mean and standard deviation of error metric in Table A2 based on five independent training trials. Additional details on data generation, training configurations, and network architecture are provided in Supplementary Information (SI) C. The code to reproduce the experiments is publicly available at https://github.com/Centrum-IntelliPhysics/TI-DeepONet-for-Stable-Long-Term-Extrapolation.git.
@@
$t_{train}=t$: Beginning time of extrapolation;   $\Delta t_{e}$: Evaluation timestep;    $T$: Final prediction time.
@@ 3.1 One-dimensional Burgers’ Equation
with periodic BCs and IC, $u(x,0)=s(x)$. We generated 2,500 realizations of $s(x)$ and the corresponding solutions, splitting them into 2,000 training and 500 test samples. To evaluate temporal extrapolation, models were trained on $t\in[0,0.5]$ and tested on the full interval $t\in[0,1]$.
In the full rollout setting, DeepONet learns a mapping $\mathcal{G}_{\boldsymbol{\theta}}\colon s(x)\mapsto u(x,t)$ using a branch network (for the IC) and a trunk network (for spatiotemporal coordinates). A similar operator learning problem is defined for FNO. In this case, the mapping is also from the initial condition $s(x)$ to the full spatiotemporal solution field $u(x,t)$. The Fourier blocks in FNO employ an FNO2D-style architecture, applying the Fourier transform across both the $x$ and $t$ coordinates. A single forward pass predicts the entire solution field. In the autoregressive setting, the models - Autoregressive DeepONet, TI-DeepONet, and TI(L)-DeepONet - the branch network takes as input the state of the system at the current timestep, and the trunk network takes only the spatial discretization. Autoregressive DeepONet predicts the solution at the next time step, while TI-DeepONet and TI(L)-DeepONet predict the time derivative of the solution at the current timestep and integrate temporal dynamics using numerical solvers. TI-DeepONet employs a Heun-type AB2/AM3 predictor–corrector scheme, defined as:
@@
Lastly, in accordance with the autoregressive DeepONet learning, autoregressive FNO also operates in an analogous manner, where the mapping is defined as $\mathcal{G}_{\boldsymbol{\theta}}\colon u(x,t)\mapsto u(x,t+1)$. An FNO1D-style architecture is utilized in this case, applying the Fourier transform only in the $x$ dimension.
@@
Figure 4 illustrates the performance of all models for a representative test sample. All models perform well within the training interval; however, the autoregressive DeepONet accumulates error early and diverges during extrapolation. The quantitative trends presented in Table 1 underscore the qualitative observations from Figure 4 as well as Figures 3(a) and (b). The autoregressive DeepONet deteriorates rapidly during extrapolation due to cumulative errors in the absence of stabilization mechanisms. While the full rollout DeepONet performs better owing to its non-recursive architecture, it fails to maintain accuracy beyond the training interval since it does not leverage the Markovian structure of time-dependent PDEs. Regarding the FNO variants, both full rollout and autoregressive FNO perform reasonably well within the training domain. However, in the extrapolation domain, both methods incur relatively higher errors, albeit lower than those of the autoregressive DeepONet. Notably, the autoregressive FNO does not exhibit the steep error growth observed in the autoregressive DeepONet during extrapolation. The full rollout FNO maintains higher errors throughout the entire temporal domain, which can be attributed to its limited capacity to capture the full spectrum of frequencies in the system, as it is trained on solution information from only half of the temporal domain.

### v3

@@ Abstract
Accurate temporal extrapolation presents a fundamental challenge for neural operators in modeling dynamical systems, where reliable predictions must extend significantly beyond the training time horizon. Conventional deep operator network (DeepONet) approaches rely on two inherently limited training paradigms: fixed-horizon rollouts, which predict complete spatiotemporal solutions while disregarding temporal causality, and autoregressive formulations, which accumulate errors through sequential predictions in time. We introduce TI-DeepONet (Time-Integrator-embedded Deep Operator Network), a framework that integrates neural operators with adaptive numerical time-stepping techniques to preserve the underlying Markovian structure of dynamical systems while substantially mitigating error propagation in extended temporal forecasting. Our approach reformulates the learning objective from direct state prediction to the approximation of instantaneous time-derivative fields, which are subsequently integrated using established numerical schemes. This design naturally supports continuous-time prediction and enables the use of higher-precision integrators at inference than those employed during training, striking a balance between computational efficiency and predictive accuracy. We further develop TI(L)-DeepONet (Learnable Time-Integrator-embedded Deep Operator Network), which incorporates learnable coefficients for intermediate stages in a multi-stage numerical integration scheme, thereby adapting to solution-specific variations and enhancing predictive fidelity. Rigorous evaluation across six canonical partial differential equations (PDEs) spanning diverse, high-dimensional, chaotic, dissipative, and dispersive dynamics shows that TI(L)-DeepONet marginally outperforms TI-DeepONet, with both frameworks achieving significant reductions in relative $L_{2}$ extrapolation error: approximately 96.3% compared to autoregressive implementations and 83.6% compared to fixed-horizon approaches. Notably, both methods maintain stable predictions over temporal domains extending to nearly twice the training interval. This research establishes a physics-aware operator learning paradigm that bridges neural approximation with numerical analysis principles, preserving the causal structure of dynamical systems while addressing a critical gap in long-term forecasting of complex physical phenomena.
@@ 1 Introduction
Time-dependent partial differential equations (PDEs) are typically solved using spatial discretization (e.g., finite element or volume) combined with time-stepping schemes such as Runge-Kutta or Adams-Bashforth. While accurate, such solvers can be prohibitively expensive for high-dimensional or many-query problems. Explicit integrators are constrained by stability conditions (e.g., CFL), while implicit schemes require solving large systems of equations at each step. Moreover, classical solvers generalize poorly: each new initial or boundary condition requires a fresh simulation.
@@
Neural operators (NOs) Lu et al. (2021); Li et al. ; Cao et al. (2024); Raonic et al. (2023); Li et al. (2020); Rahman et al. (2022); Tripura and Chakraborty (2023) offer an attractive alternative by learning mappings from input functions (e.g., initial conditions) to solutions, bypassing discretization and enabling fast inference. Models like DeepONet Lu et al. (2021); Kobayashi et al. (2024); Zhang et al. (2025); Kushwaha et al. (2024); Xu et al. (2024) and FNO Li et al. have shown promise in parametric PDE settings. However, extending NOs to time-evolving systems introduces new challenges. The standard approach is to train a model that maps the solution at time $t$ to the solution at time $t+\Delta t$, and roll this forward autoregressively Hart et al. (2023) (see Figure 1). Unfortunately, this approach suffers from severe error accumulation, as each prediction feeds into the next, compounding inaccuracies and often leading to instability. Alternatively, one can train a model to predict an entire trajectory (“full rollout”), which avoids stepwise feedback but tends to degrade in accuracy when extrapolating beyond the training time window. Crucially, most existing works Lin et al. (2023); Li et al. ; Wang and Perdikaris (2023); Lu et al. (2021) only evaluate models within the training time span, neglecting the need for reliable extrapolation, a capability that is essential for real-time engineering applications and control tasks. Moreover, existing work Xu et al. (2023); Diab and Al Kobaisi (2024); Liu and Cai (2022) on autoregressive and full rollout operator models tends to impose a discrete-time Markovian assumption, which misaligns with the underlying continuous-time dynamics of PDEs. These limitations often result in unstable or inaccurate long-term predictions. Thus, a key barrier remains: NOs struggle not only with long-horizon accuracy but also with extrapolating into unseen temporal domains, a prerequisite for trustworthy deployment.
@@
Several recent works attempt to address these limitations. For example, Diab et al. Diab and Al-Kobaisi (2025) augment DeepONet with an additional temporal branch to capture dependencies, while Michalowska et al. Michałowska et al. (2024) employ recurrent networks (RNNs, LSTMs, GRUs) atop neural operator outputs to better handle sequences. Other approaches introduce memory modules Buitrago et al. ; He et al. (2024); Koric et al. (2025); Hu et al. to compensate for the non-Markovian nature of rollout-based methods. We note that such memory-based approaches are conceptually related to the Mori-Zwanzig formalism Chorin et al. (2000), which introduces memory terms to account for unresolved dynamics; however, our approach takes a different perspective by preserving the Markovian structure inherent in the full-order system through sequential time integration. However, these strategies largely rely on generic time-series architectures rather than tackling the issue from a dynamical systems perspective.
@@
In this work, we introduce a novel approach inspired by classical numerical analysis. Our proposed TI-DeepONet framework embeds time integration directly into the operator learning process for dynamical systems. Rather than predicting future solution states directly, we train a DeepONet to approximate the instantaneous time derivative $\partial u/\partial t$. During training, this derivative operator is coupled with a Runge-Kutta numerical integrator to evolve the state forward in time. The predicted future state is then compared against the ground truth in the loss function, enabling the network to learn dynamics that respect temporal causality and numerical stability. At inference time, we leverage high-order multistep integrators (such as Adams-Bashforth/Adams-Moulton predictor-corrector schemes) with refined timesteps to reliably propagate solutions based on the learned time derivative operator. This operator learning framework which maps between function spaces rather than from function spaces to vector spaces as in standard neural networks and neural ODEs Chen et al. (2018) - provides the flexibility to modify time-stepping parameters between training and testing phases, thereby ensuring stable predictions. We further introduce TI(L)-DeepONet, an enhanced variant that incorporates learnable weighting coefficients for the intermediate stages conditioned on the system state at each timestep. This adaptation dynamically adjusts the integration process based on local solution characteristics.
@@
We demonstrate the effectiveness of our approach on six benchmark PDEs: (1) 1D Burgers’, (2) 1D Korteweg-de Vries (KdV), (3) 1D Kuramuto-Sivashinsky (KS), (4) 2D Burgers’, (5) 2D Scalar Rotation Advection-Diffusion, and (6) 3D Heat Conduction equations. We compare our methods against four standard neural operator (NO) baselines: (1) autoregressive DeepONet (DON AR), (2) full rollout DeepONet (DON FR), (3) autoregressive FNO (FNO AR), and (4) full rollout FNO (FNO FR). Across the majority of cases, TI-DeepONet and its adaptive variant TI(L)-DeepONet significantly outperform standard autoregressive and full rollout operator models, achieving stable long-term predictions and accurate extrapolation for time horizons approximately twice as long as the training interval. However, for problems involving rigid body rotational dynamics (e.g., the rotation advection-diffusion problem), FNO AR demonstrates superior performance, highlighting that problem characteristics should inform method selection.
@@ 2.2 Deep operator network (DeepONet)
Deep operator networks (DeepONet) are grounded in the universal approximation theorem for nonlinear operators Chen and Chen (1995), and are designed to learn mappings between infinite-dimensional input and output function spaces. A standard DeepONet architecture comprises two subnetworks: a branch network and a trunk network. The branch network encodes the input function $\mathbf{v}(\eta)$, evaluated at a set of fixed sensor locations $\{\eta_{1},\eta_{2},\dots,\eta_{m}\}$, producing a set of coefficient weights. The trunk network takes as input the query coordinates $\zeta=(x,y,z,t)$ and outputs the corresponding basis functions.
@@
In general, operator learning frameworks like DeepONet aim to learn a solution operator that maps any functional input, such as initial conditions (ICs), boundary conditions (BCs), or source terms, to the corresponding spatiotemporal solution field (full rollout). However, this mapping is inherently non-causal with respect to time: the operator does not explicitly enforce the Markovian nature of dynamical systems, where the future state depends on the present (or past) states. As a result, direct operator prediction over time often fails to preserve temporal coherence, particularly for long-term simulations.
@@ 2.3 Time Integrator embedded deep operator network (TI-DeepONet)
To mitigate these limitations, we propose the TI-DeepONet architecture. Rather than directly predicting the next state $u^{i+1}$, the network learns the right-hand side (RHS) of the time-dependent PDE (Eq. 1), i.e., the time derivative $u_{t}=\mathcal{F}(t,\mathbf{x},u,u_{\mathbf{x}},u_{\mathbf{xx}},\dots)$. This learned derivative is passed to a numerical time integrator to compute the next solution state from the current one. In particular, we use the RK4 scheme to advance the solution in time. For a timestep $\Delta t$, the RK4 update rule is given by, $k_{1}=\mathcal{F}(t^{i},u^{i})$, $k_{2}=\mathcal{F}\left(t^{i}+\tfrac{\Delta t}{2},u^{i}+\tfrac{\Delta t}{2}k_{1}\right)$, $k_{3}=\mathcal{F}\left(t^{i}+\tfrac{\Delta t}{2},u^{i}+\tfrac{\Delta t}{2}k_{2}\right)$, $k_{4}=\mathcal{F}\left(t^{i}+\Delta t,u^{i}+\Delta t\cdot k_{3}\right)$, and $u^{i+1}=u^{i}+\Delta t\left(\tfrac{1}{6}k_{1}+\tfrac{2}{6}k_{2}+\tfrac{2}{6}k_{3}+\tfrac{1}{6}k_{4}\right)$. This formulation leads to an end-to-end differentiable pipeline, where the loss is computed between the predicted solution $\hat{u}^{i+1}$ and the ground truth $u^{i+1}$, and gradients are backpropagated through the time integrator to update the network parameters. The loss is computed across all solution states, i.e., $N=N_{s}\times N_{t}$, where $N_{s}$ denotes the number of input realizations (samples) in the dataset and $N_{t}$ denotes the number of timesteps. However, the RK4 weighting coefficients, $\{\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}\}$, are fixed.
Boundary Conditions:
The TI-DeepONet framework is purely data-driven and does not explicitly enforce boundary conditions in the loss function. Since the network learns a one-step mapping $\mathcal{G}_{\boldsymbol{\theta}}:u^{i}\rightarrow u^{i+1}$, where both $u^{i}$ and $u^{i+1}$ are ground truth solutions that inherently satisfy the prescribed boundary conditions, the predicted outputs implicitly respect these conditions. Consequently, the framework can naturally accommodate arbitrary boundary conditions, provided the training data adheres to them.
@@ 2.4 Learnable Time Integrator embedded deep operator network (TI(L)-DeepONet)
We note that the coefficients $\boldsymbol{\alpha}=\text{NN}_{\text{RK}}(u^{i})$ are functions of the solution state $u^{i}$ rather than time $t$ directly. However, since the solution evolves temporally, the coefficients vary at each timestep, effectively introducing a time-dependence through the evolving state.
For numerical stability and interpretability, we normalize the weights using softmax activation function, such that: $\tilde{\alpha}_{j}=\frac{exp(\alpha_{j})}{\sum_{\ell=1}^{4}exp(\alpha_{\ell})},\forall j=\{1,2,3,4\}$. This adaptive scheme allows the model to adjust the influence of each intermediate slope dynamically, thereby compensating for the approximation error Lanthaler et al. (2022) of the DeepONet prediction, enabling more robust predictions in stiff, nonlinear, or highly transient regimes. For example, in regions with rapid spatial changes, the model may place greater weight on later-stage evaluations like $k_{4}$, while in smoother regions, earlier evaluations may suffice.
@@ 3 Results
To evaluate the efficacy of our framework, we consider six canonical PDEs: (1) 1D Burgers’ equation, (2) 1D KdV equation, (3) 1D KS equation, (4) 2D Burgers’ equation, (5) 2D Scalar Rotation Advection-Diffusion, and (6) 3D Heat Conduction. Model performance is assessed using the relative $L_{2}$ error, defined as $\text{error}=\frac{\|u_{\text{pred}}-u_{\text{true}}\|_{2}}{\|u_{\text{true}}\|_{2}}$, where $u_{\text{pred}}$ and $u_{\text{true}}$ denote the predicted and ground truth solutions, respectively. Table 1 and Figure 3 summarize the comparative training and extrapolation accuracy across all methods for all applications. To evaluate the statistical variability of the performance of the model, we report the mean and standard deviation of the error metrics in Table A1 based on five independent training trials. Additional details on data generation and training configurations are provided in the relevant subsections for each PDE example, with network architecture details provided in Supplementary Information (SI) B. The code to reproduce the experiments is publicly available at https://github.com/Centrum-IntelliPhysics/TI-DeepONet-for-Stable-Long-Term-Extrapolation.git.
@@
Rotating Advection Diffusion (2D) 0.27 0.008 TI(L)-DON [Ours] 0.1091 0.1332 0.1970 0.2372
TI-DON AB [Ours] 0.1091 0.1333 0.1982 0.2400
DON Full Rollout 0.1926 0.4830 0.9850 1.0562
DON Autoregressive 0.6874 0.7554 0.8925 0.9777
FNO Full Rollout 0.5093 0.5509 0.6568 0.8092
FNO Autoregressive 0.0129 0.0169 0.0286 0.0362
Heat Conduction (3D) 0.33 0.015 TI(L)-DON [Ours] 0.0142 0.0193 0.0327 0.0362
TI-DON AB [Ours] 0.0142 0.0194 0.0328 0.0364
DON Full Rollout 0.1558 0.1673 0.2058 0.2179
DON Autoregressive 0.6288 0.7421 0.9465 0.9897
$t_{train}=t$: Beginning time of extrapolation;   $\Delta t_{e}$: Evaluation timestep;    $T^{*}$: Final prediction time.
@@ 3.1 One-dimensional Burgers’ Equation
with periodic BCs and IC, $u(x,0)=s(x)$.
Data generation:
Data preparation:
The training and testing datasets are constructed differently depending on the learning strategy employed. In the full rollout setting, the goal is to train a DeepONet/FNO to learn a solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ that maps an initial condition $s(x)$ to the full spatiotemporal solution $u(x,t)$, i.e., the mapping $\mathcal{G}_{\boldsymbol{\theta}}\colon s(x)\mapsto u(x,t)$. For this purpose, we consider a set of initial conditions (input to branch network), spatiotemporal query locations (input to trunk network), and the corresponding solution fields as ground truth output. A similar operator learning problem is defined for FNO. The Fourier blocks in FNO employ an FNO 2D-style architecture, applying the Fourier transform across both the $x$ and $t$ coordinates. A single forward pass predicts the entire solution field.
The data is partitioned into training and testing sets using a train-test split ratio of 0.8, yielding $N_{\text{train}}=2000$ and $N_{\text{test}}=500$ samples. Importantly, we aim to assess the performance of various frameworks in learning the full spatiotemporal field over the entire temporal domain when given access to solutions at limited timesteps. To this end, we restrict training to only half of the temporal domain, i.e., $t_{\text{train}}\in[0,0.5]$. Consequently, the grid for the trunk network is formed by taking a meshgrid over spatiotemporal coordinates $(x,t)$, where $x\in[0,1]$ and $t\in[0,0.5]$ during training, while $t\in[0,1]$ during testing. This same meshgrid is also employed for computing the 2D Fourier Transform in FNO full rollout.
For training the autoregressive DeepONet, autoregressive FNO, TI-DeepONet, and TI(L)-DeepONet architectures, we adopt a different strategy for preparing the training and testing datasets. In autoregressive methods, future states of the system are recursively predicted using the model’s own previous outputs as inputs at each timestep. This approach inherently leads to error accumulation, as prediction errors at one timestep propagate and potentially amplify at subsequent timesteps. Understanding and mitigating this issue is a key motivation behind the TI-DeepONet architecture. To this end, the training dataset is created as follows: the input to the branch network consists of stacked solution states from $t=0$ to $t=0.5$, i.e., $u^{0}(x),u^{1}(x),\ldots,u^{50}(x)]$. Similarly, the output ground truth consists of stacked solution states from $t=0.01$ to $t=0.51$, i.e., $u^{1}(x),u^{2}(x),\ldots,u^{51}(x)]$. This arrangement enables the solution operator $\mathcal{G}_{\boldsymbol{\theta}}$ to learn the single-step mapping $u^{i}(x)\rightarrow u^{i+1}(x)$, where $u^{i}(x)$ and $u^{i+1}(x)$ represent the solution states at the $i^{\text{th}}$ and $(i+1)^{\text{th}}$ timesteps, respectively. The trunk network processes only the spatial coordinates $x\in[0,1]$. For autoregressive FNO, the inputs and outputs are distinct $u^{i}(x)$ and $u^{i+1}(x)$ solution pairs, and the Fourier transform is one-dimensional and is applied only in the $x$-dimension. Autoregressive DeepONet predicts the solution at the next time step, while TI-DeepONet and TI(L)-DeepONet predict the time derivative of the solution at the current timestep and integrate temporal dynamics using numerical solvers. TI-DeepONet employs a Heun-type AB2/AM3 predictor-corrector scheme, defined as:
@@
Training architecture:
Tables A2 and A3 present the network architecture details, including the number of layers and neurons per layer, activation functions, total epochs to convergence, batch size, and Fourier modes (where applicable). For the TI(L)-DeepONet architecture, we additionally employ an auxiliary feedforward network designed to predict the four RK4 slope coefficients. This network consists of two hidden layers with 32 neurons each, followed by an output layer with four neurons. The hidden layers use the tanh activation function, while a softmax activation is applied at the output layer to ensure the predicted coefficients sum to one. Training was performed using the Adam optimizer with an initial learning rate of $10^{-3}$, which was exponentially decayed by a factor of 0.95 every 5000 epochs. A similar learning rate scheduling strategy was employed for the auxiliary network predicting the adaptive RK4 slope coefficients. Notably for FNO AR, the learning rate schedule also employed exponential decay but by a factor of 0.96 every 2000 steps. Both training and test losses were monitored at each epoch. The final model corresponds to the set of parameters that yielded the lowest test loss over the entire training process. Among all tested architectures, TI(L)-DeepONet exhibited the fastest convergence, attributed to its capacity to adapt to the local dynamics of the solution. Following the performance results for this case, we examine the distribution of the learned RK4 slope coefficients and compare them to the reference values: $\frac{1}{6},\frac{2}{6},\frac{2}{6},\frac{1}{6}$ in a classical RK4 scheme.
@@
Figure 5: 1D Burgers’ Equation: Distribution of relative $L_{2}$ errors across all test samples for all frameworks, averaged over the entire spatiotemporal domain.
@@
Figure 4 illustrates the performance of all models for a representative test sample. All models perform well within the training interval; however, the autoregressive DeepONet (DON AR) accumulates error early and diverges during extrapolation. The quantitative trends presented in Table 1 underscore the qualitative observations from Figure 4 as well as Figures 3(a) and (b). The autoregressive DeepONet (DON AR) deteriorates rapidly during extrapolation due to cumulative errors in the absence of stabilization mechanisms. While the full rollout DeepONet (DON FR) performs better owing to its non-recursive architecture, it fails to maintain accuracy beyond the training interval since it does not leverage the Markovian structure of time-dependent PDEs. Regarding the FNO variants, both full rollout (FNO FR) and autoregressive FNO (FNO AR) perform reasonably well within the training domain. However, in the extrapolation domain, both methods incur relatively higher errors, albeit lower than those of DON AR. Notably, FNO AR does not exhibit the steep error growth observed in DON AR during extrapolation. The full rollout FNO (FNO FR) maintains higher errors throughout the entire temporal domain, which can be attributed to its limited capacity to capture the full spectrum of frequencies in the system, as it is trained on solution information from only half of the temporal domain.
Figure 6: 1D Burgers’ Equation: Distribution of the optimally learned $\alpha_{i}~\forall i\in\{1,2,3,4\}$ for the RK4 slope coefficients.